Steel can be produced using three different methods: The following table lists the average cost per ton of steel for each method.



Type of Cost Basic Continuous Electric


Process Casting Furnace



Materials 150 140 120



Labor 80 75 70



Capital 100 100 60




Energy 20 15 50





Other 45 40 25




a. Production of steel by electric furnace is a relatively new development (beginning in late 1970s) and accounts for a growing fraction of total steel sold. What is your prediction about the future production share of this method?



b. If there were a new energy crisis how would this affect steel makers choices of production methods?



c. Suppose the price of steel scrap is expected to fall significantly over the next five years. What effect would this have on the choice of production method?


1
Expert's answer
2022-02-01T14:47:26-0500

Total average cost of producing steel using basic method= sum of all cost = 395

using continous casting= 370

using electric furnace= 325

a) Since the total average cost of producing per ton of steel using electric furnace is the minimum, the production share using the same method is expected to be highest in the future.

b) When prices becomes 3 times, average cost of using energy becomes-

for basic method= 3*20= 60, total average cost becomes= 435

for continous casting=3*15=45, total average cost= 400

for electric furnace= 3*50=150, total average cost= 425

Now the producer chooses continous casting.

c) Electric furnace is chosen in part a) and it uses scap steel and when price of scap steel falls, it becomes even cheaper. So electric furnance will be chosen more.